crisulcampo:

The anode or positive terminal of an X-ray tube is at a potential of 20000 volts above the cathode or negative terminal. What work will be done on an electron by moving it between those terminals?

sillybilly123:

If you are forcing the electron toward the cathode at constant velocity, you will need to do work on the electron. If the electron is allowed to move freely toward the anode, it will accelerate gaining KE, and the electric field will be doing work on the electron. But, in either case, the **amount** of work will be: \(W = Q \cdot \Delta V = (1.602..... \times 10^{-19}) \cdot 20,000\) If you're not familiar with it, there are numerous ways to arrive at that formula. Perhaps the simplest/ most fundamental thing is to remember that a Volt is a measure of **energy per unit charge**, ie Joules per Coulomb. So, here, \(\Delta E = W = Q \cdot \Delta V\).

Bunny:

Hai der! If this question is answered, please close it. ^.^